# RecConv: Efficient Recursive Convolutions for Multi-Frequency Representations

### v2

@@ 3.1 Overall Architecture
RecNeXt-M1 | CONV | 5.2 | 0.9 | 1.4 | 361 | 384 / 3294 | 79.2 / 78.8 | ✗ | ✗ |
RecNeXt-M2 | CONV | 6.8 | 1.2 | 1.5 | 431 | 325 / 2707 | 80.3 / 80.1 | ✗ | ✗ |
@@ 4.2 Downstream Tasks
[…]
[112, 56, 28, 14] | 5.35 | 0.89 | - | - | 3221 | 75.48 | Max Unpool | |||
- | [3, 3, 3, 3] | 4.80 | 2.40 | 2.07 | 13.2 | 1553 | 76.35 | - | ||
[7, 7, 7, 7] | 4.96 | 2.55 | 2.20 | 13.2 | 1224 | 77.42 | ||||
[11, 11, 11, 11] | 5.26 | 2.82 | 2.27 | 13.8 | 880 | 78.00 | ||||
[…]
@@ 4.3 Ablation Studies
Upsampling operation. Table 5 demonstrates the effectiveness of substituting bilinear upsampling with transposed depthwise convolution for images, While this replacement slightly increases computational complexity and parameter count, it potentially enhances feature fidelity during upsampling, resulting in a modest increase in top-1 accuracy from 78.03% to 78.24% with kernels. Additionally, it reduces latency from 2.84ms to 2.63ms and boosts throughput from 300img/s to 580img/s. However, smaller kernel sizes (e.g., and ) lead to performance degradation, with accuracy dropping from 78.11% to 77.87% and from 77.72% to 77.26%, respectively.
RecConv Variants. We present RecConv variants (Figure 3) incorporating recursive decomposition along both spatial and channel dimensions. As summarized in Table 5, the variant with group convolutions reduces parameter from 5.81M to 5.76M, FLOPs from 2.75G to 2.70G, and increases throughput from 576 to 636, but results in a 0.33% drop in accuracy (from 78.24% to 77.91%). Meanwhile, the latency increases significantly due to the lack of adaptation of the transposed group convolution on the NPU. The variant using channel-wise concatenation further reduces parameters by 0.4M and FLOPs by 0.1G, lowers latency by 0.3ms, and increases throughput by 200, but causes a more significant accuracy drop of 1.4%.
@@
5 Conclusions
Supplementary Material
6 RecConv with Linear Attention
@@
A5 | 25.7 | 4.7 | 5.6 | 1376 | 721 | 83.1 / 83.5 |
As detailed in Table 7, the RecNeXt models offer a compelling trade-off between accuracy, model size, and hardware-specific efficiency, achieving up to 83.1% (83.5% with knowledge distillation) Top-1 accuracy on ImageNet-1K with models from 2.5M to 25.7M parameters. The architecture provides two specialized paths: the M-series for accelerator-centric deployment with NPU latencies as low as 1.0ms, and the A-series for GPU-bound scenarios, where its linear attention and nearest interpolation design achieves higher throughput and better accuracy.
@@
Table 8 presents a comparative analysis of our model families, evaluated across three distinct computational budgets on object detection, instance segmentation, and semantic segmentation tasks. At the most compact scale, the A3 model demonstrates superiority over M3 across all evaluated metrics while also being computationally efficient. As model complexity increases, a nuanced performance trade-off emerges. The M-series generally excels in object detection and instance segmentation. In contrast, the A-series consistently shows a competitive edge in semantic segmentation, with A3 achieving a notable +0.9 mIoU over M3 and A5 surpassing M5 by +0.5 mIoU.
7 Shape Bias Analysis
@@
8 More Comparisons
@@
RecNeXt-B | 19.3 | 1.1 | 2.5 | 290 | 4385 / 4232 | 80.3 / 81.4 |
We evaluate “T”, “S”, “B” variants on ImageNet-1K (Table 10), comparing them against the SOTA efficient architecture, LSNet [71]. The NPU latency is measured on an iPhone 13 with models compiled by Core ML Tools. The CPU latency is accessed on a Quad-core ARM Cortex-A57 processor in ONNX format. And the throughput is tested on a Nvidia RTX3090 with maximum power-of-two batch size that fits in memory. “” denotes hard knowledge distillation using the RegNetY-16GF [51] as the teacher model.
The smallest model, RecNeXt-T, achieves a Top-1 accuracy of 75.1%, which is a 0.2% improvement over LSNet-T, with a negligible increase in parameters and identical FLOPs. With knowledge distillation, RecNeXt-T reaches 76.6% accuracy, surpassing LSNet-T by 0.5%. The model demonstrates impressive hardware efficiency, with an NPU latency of 1.8ms, an ARM CPU latency of 109ms, and a GPU throughput of 14,181 images per second. RecNeXt-S achieves 78.3% Top-1 accuracy, outperforming LSNet-S by 0.5%, while having less parameters. The distilled version of RecNeXt-S widens this gap, reaching 79.6% accuracy. RecNeXt-B matches the 80.3% Top-1 accuracy of LSNet-B with fewer FLOPs and higher throughput. Across all model sizes, our architecture consistently delivers a competitive trade-off between accuracy, latency, and model size compared to LSNet. Table 11 presents the detailed architecture, where the “RecAttn Block” denotes partial channel operation [5, 76] of the RecConv block with linear attention.
@@
1 | RecAttn Block | channels | 64 | 128 | 128 | |
@@
2 | RecAttn Block | channels | 128 | 256 | 256 | |
@@
3 | RecAttn Block | channels | 256 | 384 | 384 | |
@@
4 | RecAttn Block | channels | 512 | 512 | 512 | |
blocks | 10 | 10 | 12 |
@@ 9 Training Details
For ImageNet-1K classification benchmark, we adopt the similar training recipe as RepViT [69] and LSNet [71]. All models are trained from scratch for 300 epochs on 224224 resolution. The initial learning rate is set to , and the total batch size is set to 1024. Table 12 provides the detailed training hyper-parameters.

### v3

@@ 3.1 Overall Architecture
| RecNeXt-M1 | CONV | 5.2 | 0.9 | 1.4 | 361 | 384 / 3303 | 79.2 / 78.8 | ✗ | ✗ |
| RecNeXt-M2 | CONV | 6.8 | 1.2 | 1.5 | 431 | 325 / 2724 | 80.3 / 80.1 | ✗ | ✗ |
@@ 4.2 Downstream Tasks
| [112, 56, 28, 14] | 5.72 | 0.99 | 1.46 | 17.2 | 2077 | 75.93 | Nearest Interpolate | |||
| [112, 56, 28, 14] | 5.55 | 0.91 | 1.43 | 14.6 | 2677 | 75.71 | ||||
| [112, 56, 28, 14] | 5.35 | 0.89 | - | - | 2770 | 75.48 | Max Unpool | |||
| - | [3, 3, 3, 3] | 4.80 | 2.40 | 2.07 | 13.2 | 1575 | 76.35 | - | ||
| [7, 7, 7, 7] | 4.96 | 2.55 | 2.20 | 13.2 | 1239 | 77.42 | ||||
| [11, 11, 11, 11] | 5.26 | 2.82 | 2.27 | 13.8 | 889 | 78.00 | ||||
| [4, 3, 2, 1] | [48, 24, 12, 6] | 4.90 | 2.44 | 2.71 | 20.6 | 320 | 77.72 | Bilinear Interpolate | ||
@@ 4.3 Ablation Studies
Upsampling operation. Table 5 demonstrates the effectiveness of substituting bilinear upsampling with transposed depthwise convolution for images, While this replacement slightly increases computational complexity and parameter count, it potentially enhances feature fidelity during upsampling, resulting in a modest increase in top-1 accuracy from 78.03% to 78.24% with kernels. Additionally, it reduces latency from 2.84ms to 2.63ms and boosts throughput from 300img/s to 500img/s. However, smaller kernel sizes (e.g., and ) lead to performance degradation, with accuracy dropping from 78.11% to 77.87% and from 77.72% to 77.26%, respectively.
RecConv Variants. We present RecConv variants incorporating recursive decomposition along both spatial and channel dimensions. As summarized in Table 5, the variant with group convolutions reduces parameter from 5.81M to 5.76M, FLOPs from 2.75G to 2.70G, and increases throughput from 502 to 563, but results in a 0.33% drop in accuracy (from 78.24% to 77.91%). Meanwhile, the latency increases significantly due to the lack of adaptation of the transposed group convolution on the NPU. The variant using channel-wise concatenation further reduces parameters by 0.4M and FLOPs by 0.1G, lowers latency by 0.3ms, and increases throughput by 240, but causes a more significant accuracy drop of 1.4%.
RecConv Aggregation. We reformulate RecConv’s up-sampling aggregation (Eq. 5) to establish its connection to recurrent architectures like RNNs [61] and SSMs [18, 17]. This perspective views the downsampled feature maps across decomposition levels as a sequential input to a recurrent model (Eq. 6). Where represents the weight for the convolution of level , while and are the weights of the hidden state and input, respectively. denotes the spatial upsampling operation. This modification slightly boosts accuracy but increases latency and FLOPs (detailed in Algorithm 2 and Table 5).
@@
4.4 RecConv with Linear Attention
@@
| A5 | 25.7 | 4.7 | 5.6 | 1376 | 733 | 83.1 / 83.5 |
RecConv is designed with great flexibility, enabling its convolutional modules to be easily replaced by alternative operations like linear attention [33, 21]. As detailed in Table 7, the RecNeXt models offer a compelling trade-off between accuracy, model size, and hardware-specific efficiency, achieving up to 83.1% (83.5% with knowledge distillation) Top-1 accuracy on ImageNet-1K with models from 2.5M to 25.7M parameters. The architecture provides two specialized paths: the M-series for accelerator-centric deployment with NPU latencies as low as 1.0ms, and the A-series for GPU-bound scenarios, where its linear attention and nearest interpolation design achieves higher throughput and better accuracy.
@@
Table 8 presents a comparative analysis of our model families, evaluated across three distinct computational budgets on object detection, instance segmentation, and semantic segmentation tasks. At the most compact scale, the A3 model demonstrates superiority over M3 across all evaluated metrics while has fewer FLOPs. As model complexity increases, a nuanced performance trade-off emerges. The M-series generally excels in object detection and instance segmentation. In contrast, the A-series consistently shows a competitive edge in semantic segmentation, with A3 achieving a notable +0.9 mIoU over M3 and A5 surpassing M5 by +0.5 mIoU.
4.5 Shape Bias Analysis
@@
4.6 More Comparisons
To further validate the effectiveness of RecConv, we compare it against the SOTA efficient architecture, LSNet [78] on ImageNet-1K. We adopt the architecture design of LSNet, but replace the token mixer with partial channel [5, 83] MetaNeXt [47, 82] style. With shared channel operation in the final stage to further reduce computational cost and minimize feature redundancy, the detailed architecture is shown in Figure 4. Where the “RecLinear Attention” denotes using linear attention in the RecConv block.
@@ 4.6 More Comparisons
| RecNeXt-B | 19.2 | 1.1 | 2.5 | 296 | 4472 / 4353 | 80.3 / 81.5 |
We evaluate “T”, “S”, “B” variants on ImageNet-1K as shown in Table 10. The NPU latency is measured on an iPhone 13 with models compiled by Core ML Tools. The CPU latency is accessed on a Quad-core ARM Cortex-A57 processor in ONNX format. And the throughput is tested on a Nvidia RTX3090 with maximum power-of-two batch size that fits in memory. “” denotes hard knowledge distillation using the RegNetY-16GF [56] as the teacher model.
The smallest model, RecNeXt-T, achieves a Top-1 accuracy of 75.2%, which is a 0.3% improvement over LSNet-T, with a negligible increase in parameters and identical FLOPs. With knowledge distillation, RecNeXt-T reaches 76.8% accuracy, surpassing LSNet-T by 0.7%. The model demonstrates impressive hardware efficiency, with an NPU latency of 1.8ms, an ARM CPU latency of 105ms, and a GPU throughput of 13,957 images per second. RecNeXt-S achieves 78.3% Top-1 accuracy, outperforming LSNet-S by 0.5%, while having less parameters. The distilled version of RecNeXt-S remains this gap, reaching 79.5% accuracy. RecNeXt-B matches the 80.3% Top-1 accuracy of LSNet-B with fewer FLOPs and higher throughput. Across all model sizes, our architecture consistently delivers a competitive trade-off between accuracy, latency, and model size compared to LSNet. Table 11 presents the detailed architecture, where the “RecAttn Block” denotes partial channel operation [5, 83] of the RecConv block with linear attention.
5 Conclusions
Supplementary Material
6 RecConv Aggregation
Algorithm 2 shows the implementation of the RecConv aggregation like SSM in a PyTorch-like style. this approach treats the sequence of downsampled feature maps from each decomposition level as the input to a recurrent model.
7 Grad-CAM Visualization
We visualize class activation maps (CAM) using Grad-CAM [62] with the TorchCAM Toolbox [14]. As shown in Figure 5, RecNeXt achieves a better trade-off between local and global features compared to other efficient models.
@@ 7 Grad-CAM Visualization
| 1 | RecAttn Block | channels | 64 | 128 | 128 | |
@@
| 2 | RecAttn Block | channels | 128 | 256 | 256 | |
@@
| 3 | RecAttn Block | channels | 256 | 384 | 384 | |
@@
| 4 | RecAttn Block | channels | 512 | 512 | 512 | |
| blocks | 10 | 10 | 12 | |||
@@ 8 Training Details
We adopt the training recipe form RepViT [76] and LSNet [78] for ImageNet-1K classification. All models are trained from scratch for 300 epochs using 224224 resolution input. The initial learning rate is set to , and the total batch size is set to 1024. DropPath [36] is applied to the larger variants for regularization without knowledge distillation. Table 12 provides the detailed training hyper-parameters.